Construct a depreciation. Schedule for the following problems.

1.) The Dwight laundry purchased new washing machines and dryers for ₱570,000.Freight charges were ₱47,000 and installation amounted to ₱50,000.The machines are expected to last five years and have a salvage value or scrap value ,₱20,000. If Dwight elects to use the straight-line method of depreciation, prepare a depreciation schedule for these machines .


1
Expert's answer
2021-05-18T14:14:53-0400

Total cost = cost + shipping charges + installation charges

= 570,000 + 47,000 + 50,000

= 667,000

Total depreciation = Total cost – Salvage value

= 667,000 – 20,000

= 647,000

Annual depreciation "= \frac{Total \; depreciation}{Number \;of\; useful \;years}"

"= \frac{647,000}{5}"

= 129,400

Original book value = 667,000

At the end of the 1st year

Annual depreciation = 129,400

Accumulated depreciation = 129,400

Book value (at the end of the 1st year)

= Original book value – Annual depreciation

= 667,000 – 129,400

= 537,600


At the end of the 2nd year

Annual depreciation = 129,400

Accumulated depreciation = Annual depreciation "\times 2"

= 129,400 "\times 2"

= 258,800

Book value (at the end of the 2nd year)

= Book value (at the end of the 1st year) – Annual depreciation

= 537,600 – 129,400

= 408,200


At the end of the 3rd year

Annual depreciation = 129,400

Accumulated depreciation = Annual depreciation "\times 3"

= 129,400 "\times 3"

= 388,200

Book value (at the end of the 3rd year)

= Book value (at the end of the 2nd year) – Annual depreciation

= 408,200 – 129,400

= 278,800


At the end of the 4th year

Annual depreciation = 129,400

Accumulated depreciation = Annual depreciation "\times 4"

= 129,400 "\times 4"

= 517,600

Book value (at the end of the 4th year)

= Book value (at the end of the 3nd year) – Annual depreciation

= 278,800 – 129,400

= 149,400


At the end of the 5th year

Annual depreciation = 129,400

Accumulated depreciation = Annual depreciation "\times 5"

= 129,400 "\times 5"

= 647,000

Book value (at the end of the 5th year)

= Book value (at the end of the 4nd year) – Annual depreciation

= 149,400 – 129,400

= 20,000